For a particular reaction at 134.8 °C, ΔG = -1319.75 kJ/mol, and ΔS = 863.35 J/(mol·K). Calculate ΔG for this reaction at -79.1 °C.
Answer #1

Given, \Delta G = -1319.75 kJ/mol = -1319.75 *10^3 J/mol

\DeltaS = 863.35 J/mol K and T1 = 134.8 oC = 273.15+134.8 = 407.95K

We know, \Delta G = \Delta H - T​1\DeltaS , then \Delta H = \Delta G + T\DeltaS =   -1319.75 *10^3 + (407.95*863.35) = -967.546*10^3 J/mol = -967.546kJ/mol

Now, at T2 = -79.1 oC = -79.1+273.15 = 194.05 K

\DeltaG = \Delta H - T​1\DeltaS = (-967.546*10^3) - (194.05*863.35) = -1135.079 *10^3 J/mol = -1135.079 kJ/mol